OpenStudy (anonymous):

18r^4s^5t^6 (numerator) -3r^2st^3 (denominator) equals what?

OpenStudy (anonymous):

subtract exponents when dividing like terms

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

for example: x^6/x^4=x^2 6-4=2

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[\frac{18 r^4 s^5 t^6}{-3 r^2 s t^3}=-6 r^2 s^4 t^3 \]